OpenStudy (anonymous):

4.The hypotenuse of a 30°- 60°- 90° right triangle measures 22 cm. What are the lengths of the longer leg and shorter leg?

OpenStudy (anonymous):

yahoo you know uniform triangle?

OpenStudy (anonymous):

no but if u want help

OpenStudy (anonymous):

I think shorter leg=22/2=11 and longer leg = 22^2-11^2